Example question: infuse vancomycin 500mg in 40 mL of DsW infusion bag at 10 mcg per kg per minute. The client weighs 176 lb. what will the pump setting be for continuous drip find the rate and VTBI how many hours will the IV bag last How many minutes will the IV bag last
Answer #1

1) The contnuous drip is a controlled method of IV adminstration of drugs , fluids or nutrients given without interruption instead of bolus and the pump setting for continuous drip is small volume pumps infuse medicines , insulin etc and small volume pumps usually use a computer controlled motor turning screw that pushes the plunger on a syringe, the rate (ml/hr) =volume/time so the aswer is 4ml( 40ml/10), and VTBI (Volume to be infused) it is 40ml, 2) first the mcg/kg/min is converted iinto mg/hr , 10*80/1000*60=48 hrs, IV dosage formula= 48/500*40=3.84ml/hr, time=volume(ml)/rate(ml/hr) so the hours taken the IV bag lasts is 10hours(10.4). 3) it takes 600 minutes(10 hours) for the IV bag to last.